OpenStudy (ell):

Solve each equation by factoring: b^2 + 5b - 35 = 3b

OpenStudy (tyteen4a03):

Firstly move 3b to the other side so you get the standard quadratic equation, which looks like Ax^2 + Bx + C = 0.

OpenStudy (ell):

b^2 + 2b -35 = 0

OpenStudy (anonymous):

once you do that then find 2 numbers that multyply to -35 and add up to (5-3)

OpenStudy (anonymous):

yes now what mulitplies to -35 and adds up to 2?

OpenStudy (ell):

Thanxs! haha(: give me a sec

OpenStudy (ell):

haha(: idk!

OpenStudy (ell):

Thanx everyone! GBU!! Much appreciated! <3

OpenStudy (anonymous):

-5*7=35 7-5

OpenStudy (anonymous):

so b^2 + 2b -35 = 0 = (b-5)*(b+7) so solve (b-5)*(b+7)=0

OpenStudy (ell):

Thank you so much!